# How to warm-start your unfolding network

### v1

@@ I Introduction
Output : solution $\hat{x}$ of (1)
1 Initialize $z_{0}^{1}\in\mathbb{R}^{N}$, $z_{0}^{2}\in\mathbb{R}^{m}$;
2 $u_{0}^{1}=z_{0}^{1},\,u_{0}^{2}=z_{0}^{2}$;
@@ II Main results
Model-based CS: CS pertains to reconstructing $x\in\mathbb{R}^{n}$ from $y=Ax+e\in\mathbb{R}^{m}$, $m<n$, with $A\in\mathbb{R}^{m\times n}$ and $e\in\mathbb{R}^{m}$, $\|e\|_{2}\leq\varepsilon$. The data $x$ is assumed to be sparse after the application of a known transform $W$, to ensure exact/approximate reconstruction; in this paper, we consider the case of a redundant $W\in\mathbb{R}^{N\times n}$. In terms of optimization, this is translated into minimizing a smooth, sparsity-inducing objective function satisfying a data-fitting term. In other words, a common approach [tfocs] for formulating the CS problem is:
@@
with $x_{0}\in\mathbb{R}^{n}$ being an initial guess on $x$, and $\mu>0$. A state-of-the-art algorithm solving (1) is Algorithm 1. The latter takes as input $y$, introduces dual variables $z^{1}\in\mathbb{R}^{N}$, $z^{2}\in\mathbb{R}^{m}$ and step size sequences $\{t_{k}^{1}\},\,\{t_{k}^{2}\},\{\theta_{k}\}>0$, and produces an iterative scheme, which after some iterations outputs $\hat{x}\approx x$.
Due to the appearance of the term $(\mu/2)\|x-x_{0}\|_{2}^{2}$ in (1), the performance of the solver depends on the choice of $x_{0}$; to alleviate this effect, one can use continuation. The latter is an algorithmic method deployed to increase the solver’s performance and hinges upon iteratively solving (1) with better and better guesses on $x_{0}$. In other words, continuation solves a sequence of similar but easier problems, using the results of each “subproblem” to “warm start”, namely, initialize the next one. The combination of Algorithm 1 with continuation results in Algorithm 2. For the remainder of the paper, we focus on updating $x_{0}$ and keep a fixed value of $\mu$. Of particular interest is the fact that continuation is quite efficient when coupled with Algorithm 1 and boosts its performance, in terms of reconstruction error [tfocs].
@@
with $v\in\mathbb{R}^{(2N+2m)\times 1}$ being an auxiliary variable and $L$ the total number of layers. Formally, $y$ is passed through all subsequent layers, but for the sake of readability, we only write the dependence on $v$ here. The matrices $\{D_{k},\Theta_{k}\}_{k\geq 1}\in\mathbb{R}^{(2N+2m)\times(2N+2m)}$ and the nonlinear function $\sigma(\cdot)$ depend on the step sizes and the optimization variables of Algorithm 1 (see [deconet] for more implementation details). By composing the outputs layer after layer, DECONET implements a decoder $h^{W}:\mathbb{R}^{m}\mapsto\mathbb{R}^{n}$ for CS, such that $h^{W}(y):=h(y)=\hat{x}\approx x$. The learnable $W$ is shared across all layers, and due to its redundancy, it renders DECONET as an overparameterized DUN. Since both DECONET and Algorithm 1 solve (1), we can replace the latter with the former inside the continuation loop of Algorithm 2. This tantamounts to a new unfolded architecture, where we combine the continuation loop with the DUN, presented in Algorithm 3. We coin this hybrid model-based unfolded architecture continued decoder (C-DEC) and write $g_{x_{0}}(y)$ for the decoder it implements, to distinguish it from the decoder implemented by DECONET.
Given a training sequence of $s$ pair-samples, we aim to measure the difference between $x_{i}$ and $\hat{x}_{i}^{\star}=g_{x_{0}}(y_{i})$, $i=1,\dots,s$. To that end, we choose the log-cosh loss function [logcosh]:
$\mathcal{L}=\frac{1}{s}\sum_{i=1}^{s}\log\cosh(g_{x_{0}}(y_{i})-x_{i}),$ (4)
as opposed to DECONET, which is trained and tested with the mean-squared error (MSE) [deconet]. Our choice of log-cosh [statlogcosh] is attributed to its behavior as the MSE close to the origin, and as the mean-absolute error [superres] far from the origin. Hence, log-cosh incorporates properties of both loss functions, which are standard metrics for DUNs’ performance. Based on this aspect and the success of continuation when coupled with the iterative solver for CS, we surmise that C-DEC will demonstrate an improved performance compared to its vanilla counterpart. In Sec. III, we provide empirical evidence, highlighting the superiority of the proposed ensemble framework.
Input : Solution $\hat{x}$ of (1)
Output : Warm-started solution $\hat{x}^{\star}$ of (1)
1 $\mu_{0}>0$;
@@ 4       $x_{0}^{j+1}\leftarrow\hat{x}^{j}+\frac{j}{j+3}(\hat{x}^{j+1}-\hat{x}^{j})$;
5       (optional) increase or decrease of $\mu_{j}$;
Algorithm 2 Continuation on Alg. 1 [tfocs, Listing 11]
Input : Solution $\hat{x}$ of (1)
Output : Warm-started solution $\hat{x}^{\star}$ of (1)
@@ III Experiments
Settings: We train and test C-DEC on MNIST (60000 training and 10000 test $28\times 28$ image examples) and CIFAR10 (50000 training and 10000 test $32\times 32$ coloured image examples). For the latter, we transform the images into grayscale ones. For both datasets, we consider the vectorized form of the images. For the CS measurements, we fix their number $m=0.25\times n$, use a Gaussian $\tilde{A}:=A/\sqrt{m}\in\mathbb{R}^{m\times n}$, and add zero-mean Gaussian noise $e$ with standard deviation $\mathrm{std}=10^{-4}$ to the measurements $y$, so that $y=\tilde{A}x+e$. We set $\varepsilon=\|y-\tilde{A}x\|_{2}$ and $x_{0}=A^{T}y$, which are standard algorithmic setups for Algorithm 1. We investigate C-DEC with varying number of layers and continuation steps. For the learnable sparsifying transform $W\in\mathbb{R}^{N\times n}$ we consider two instances, with $N=10\cdot n$ and $N=50\cdot n$, both initialized with a Beta distribution [dyingrelu]. The rest of C-DEC’s hyperparameters, e.g., $\mu$, $t^{1},t^{2},\theta$, are inherited by DECONET and thus set according to [deconet]. All models are implemented in PyTorch [pytorch] and trained using Adam algorithm [adam], with batch size $128$ and learning rates $\eta=10^{-3}$ for MNIST and $\eta=10^{-4}$ for CIFAR10. For our experiments, we report the test log-cosh loss $\mathcal{L}_{test}$ – which is essentially the log-cosh loss evaluated on a set of $d$ test data not used during training – and the empirical generalization error (EGE)
$\mathcal{L}_{gen}=|\mathcal{L}_{test}-\mathcal{L}|,$ (5)
with $\mathcal{L}$ given in (4). We train C-DEC, on all datasets, employing an early stopping technique [earlystop] with respect to (5). We repeat all the experiments at least 10 times and average the results over the runs. Since our main objective is to showcase how the model-based continuation ripples out to the behavior of the unfolded network, we only employ DECONET as a baseline and leave other DUNs as future work, especially since [deconet] has demonstrated that DECONET outperforms two other state-of-the-art DUNs. For the loss landscapes, we follow the standard framework of [lossvis] and perturb the learned $W$ in two random directions, with scalars $Alpha$ and $Beta$ representing the extent of the perturbation in each direction.
@@
Results & discussion: We examine the reconstruction and generalization performance of 10- and 15-layer C-DEC, with $W\in\mathbb{R}^{10n\times n}$ and varying number of continuation steps, on both datasets, and illustrate the findings in Fig. 1. From a model-based viewpoint, the decays in test errors seem reasonable: as the number of continuation steps (and layers) increases, the solver’s reconstruction ability is improved [tfocs]. This behavior conforms with our intuition presented in Sec. II. Interestingly, we find that the number of continuation steps is mildly reversely proportional to the generalization error. We surmise that this is due to the updated architecture of C-DEC compared to DECONET, although further mathematical exploration is needed. Additionally, we compare C-DEC and DECONET, both with $W\in\mathbb{R}^{50n\times n}$, for 10 and 50 layers, on all datasets, and present the findings in Table I. Both the test and generalization errors are always lower for our proposed DUN, consistently for both datasets. Based on the results, we obtain a two-fold insight: a) we confirm our model-based intuition that continuation boosts the solver’s performance, even in the unfolding regime b) we justify our choice of the log-cosh loss as an adequate metric for training and evaluating the proposed DUN, compared to MSE, which is deployed for DECONET. Overall, results highlight that continuation is a favorable component of C-DEC’s architecture, with the log-cosh loss being a better choice over standard metrics, e.g., the MSE, for measuring DUNs’ performance.
Finally, we examine how continuation ripples out to C-DEC’s loss landscapes. We compare a 5-layer C-DEC without continuation – which is architecturally equivalent to DECONET – and a 5-layer C-DEC with 5 steps of continuation. We examine both DUNs with $W\in\mathbb{R}^{10n\times n}$ and $W\in\mathbb{R}^{50n\times n}$, report the results in Fig. 2. The illustrations demonstrate a very intriguing phenomenon. Firstly, we notice that for both datasets, DECONET (equivalent to C-DEC with 1-step continuation) suffers from a highly nonsmooth loss landscape, with many peaks and valleys. On the other hand, C-DEC with 5 steps of continuation enjoys a much smoother loss landscape, with the effect being highly noticeable for the CIFAR10 images, as depicted in Fig. 2b. Then, for the MNIST dataset illustrated in Fig. 2a, we notice that overparameterization also smooths out the landscape, as $N$ increases from $10n$ to $50n$. Although this effect is negligible for DECONET, it is more noticeable for C-DEC, showcasing the potential of coupling continuation with overparameterization in the context of DUNs. All in all, results indicate that continuation (and overparameterization) has a strong positive effect on the optimization process of DUNs and could spark an interesting mathematical theory regarding DUNs’ generalization.

### v2

@@ I Introduction
Output : solution $\hat{x}\in\mathbb{R}^{n}$ of (1)
1 Initialize $x_{0}\in\mathbb{R}^{n}$, $z_{0}^{1}\in\mathbb{R}^{N}$, $z_{0}^{2}\in\mathbb{R}^{m}$;
2 $\theta_{0}\leftarrow 1$, $u_{0}^{1}\leftarrow z_{0}^{1},\,u_{0}^{2}\leftarrow z_{0}^{2}$;
@@ II Main results
Model-based CS: CS pertains to reconstructing $x\in\mathbb{R}^{n}$ from $y=Ax+e\in\mathbb{R}^{m}$, $m<n$, with $A\in\mathbb{R}^{m\times n}$ and $e\in\mathbb{R}^{m}$, $\|e\|_{2}\leq\varepsilon$. The data $x$ is assumed to be sparse after the application of a known transform $W$, to ensure exact/approximate reconstruction; in this paper, we consider the case of a redundant $W\in\mathbb{R}^{N\times n}$. In optimization terms, this is translated into minimizing a smooth, sparsity-inducing objective function satisfying a data-fitting term. In other words, a common approach [tfocs] for formulating the CS problem is:
@@
with $x_{0}\in\mathbb{R}^{n}$ being an initial guess on $x$, and $\mu>0$. A state-of-the-art algorithm solving (1) is Algorithm 1. The latter takes as input $y$, introduces dual variables $z^{1}\in\mathbb{R}^{N}$, $z^{2}\in\mathbb{R}^{m}$ and step sizes $t^{1},\,t^{2},\,\theta$, and produces an iterative scheme, which after some iterations outputs $\hat{x}\approx x$.
Due to the appearance of the term $(\mu/2)\|x-x_{0}\|_{2}^{2}$ in (1), the performance of the solver depends on the choices of $x_{0}$ (and $\mu$); to alleviate this effect, one can use continuation. The latter is an algorithmic method deployed to increase the solver’s performance and hinges upon iteratively solving (1) with better and better guesses on $x_{0}$ (and $\mu$). In other words, continuation solves a sequence of similar but easier problems, using the results of each “subproblem” to “warm start”, i.e., initialize the next one. The combination of Algorithm 1 with continuation results in Algorithm 2, presented here slightly differently from the one in [tfocs], to satisfy our paper’s formulation. We focus on updating $x_{0}$, and keep $\mu$ fixed, since updating the latter is optional [tfocs]. Of particular interest is the fact that continuation is quite efficient when coupled with Algorithm 1 and boosts its performance, in terms of reconstruction error [tfocs].
@@
with $v\in\mathbb{R}^{(2N+2m)\times 1}$ being an auxiliary variable and $L$ the total number of layers. Formally, $y$ is passed through all subsequent layers, but for the sake of readability, we only write the dependence on $v$ here. The matrices $\{D_{k},\Theta_{k}\}_{k\geq 1}\in\mathbb{R}^{(2N+2m)\times(2N+2m)}$ and the nonlinear function $\sigma(\cdot)$ depend on the step sizes and the optimization variables of Algorithm 1 (see [deconet] for more implementation details). By composing the outputs layer after layer, DECONET implements a decoder $h^{W}:\mathbb{R}^{m}\mapsto\mathbb{R}^{n}$ for CS, such that $h^{W}(y):=h(y)=\hat{x}\approx x$. The learnable $W$ is shared across all layers, and due to its redundancy, it renders DECONET as an overparameterized DUN. Since both DECONET and Algorithm 1 solve (1), we can replace the latter with the former inside the continuation loop of Algorithm 2. This tantamounts to a new unfolded architecture, presented in Algorithm 3, where we combine the continuation loop with DECONET. We coin this hybrid model-based unfolded architecture continued decoder (C-DEC) and write $g^{W}_{x_{0}}(y)$ for the decoder it implements, to distinguish it from the decoder implemented by DECONET.
Given a training sequence of $s$ pair-samples, we aim to measure the difference between $x_{i}$ and $\hat{x}_{i}^{\star}=g^{W}_{x_{0}}(y_{i})$, $i=1,\dots,s$. To that end, we choose the log-cosh loss function [logcosh]:
$\mathcal{L}_{\mathrm{train}}=\frac{1}{s}\sum_{i=1}^{s}\log\cosh(g^{W}_{x_{0}}(y_{i})-x_{i}),$ (4)
as opposed to DECONET, which is trained and tested with the mean-squared error (MSE) [deconet]. Our choice of log-cosh [statlogcosh] is attributed to its behavior as the MSE close to the origin, and as the mean-absolute error [superres] far from the origin. Hence, log-cosh incorporates properties of both loss functions, which are standard metrics for DUNs’ performance. Based on this aspect and the success of continuation when coupled with the iterative solver for CS, we believe that C-DEC will demonstrate an improved performance compared to its vanilla counterpart. In Sec. III, we provide empirical evidence, confirming the superiority of the proposed ensemble framework.
Input : $y\in\mathbb{R}^{m}$
Output : Warm-started solution $\hat{x}^{\star}\in\mathbb{R}^{n}$ of (1)
1 Initialize $x_{0}^{0}\in\mathbb{R}^{n}$, $\mu_{0}>0$;
@@ 4       $x_{0}^{j+1}\leftarrow\hat{x}^{j}+\frac{j}{j+3}(\hat{x}^{j+1}-\hat{x}^{j})$;
5       (optional) linear increase or decrease of $\mu_{j}$;
Algorithm 2 Alg. 1 with continuation [tfocs, Listing 11]
Input : $y\in\mathbb{R}^{m}$
Output : Warm-started solution $\hat{x}^{\star}\in\mathbb{R}^{n}$ of (1)
1 Initialize $x_{0}^{0}\in\mathbb{R}^{n}$;
@@ III Experiments
Settings: We train and test C-DEC on MNIST (60000 training and 10000 test $28\times 28$ image examples) and CIFAR10 (50000 training and 10000 test $32\times 32$ coloured image examples). We transform the CIFAR10 images into grayscale ones. For both datasets, we consider the vectorized form of the images. We fix $m=0.25\cdot n$, use a Gaussian $\tilde{A}:=A/\sqrt{m}\in\mathbb{R}^{m\times n}$, and add zero-mean Gaussian noise $e$ with standard deviation $\mathrm{std}=10^{-4}$ to the measurements $y$, so that $y=\tilde{A}x+e$. We set $\varepsilon=\|y-\tilde{A}x\|_{2}$ and $x_{0}=A^{T}y$, which are standard algorithmic setups for Algorithm 1. We investigate C-DEC with varying number of layers and continuation steps. For the learnable sparsifying transform $W\in\mathbb{R}^{N\times n}$ we consider two instances, with $N=10\cdot n$ and $N=50\cdot n$, both initialized with a Beta distribution [dyingrelu]. The rest of C-DEC’s hyperparameters, e.g., $\mu$, $t^{1},t^{2},\theta$, are inherited by DECONET and thus set according to [deconet]. All models are implemented in PyTorch [pytorch] and trained using Adam algorithm [adam], with batch size $128$ and learning rates $\eta=10^{-3}$ for MNIST and $\eta=10^{-4}$ for CIFAR10. For our experiments, we report the test log-cosh loss $\mathcal{L}_{\mathrm{test}}$ – which is essentially the log-cosh loss evaluated on a set of $d$ test data not used during training – and the empirical generalization error (EGE) $\mathcal{L}_{\mathrm{gen}}=|\mathcal{L}_{\mathrm{test}}-\mathcal{L}_{\mathrm{train}}|$, with $\mathcal{L}_{\mathrm{train}}$ as in (4). We train C-DEC, on all datasets, employing an early stopping technique [earlystop] with respect to $\mathcal{L}_{\mathrm{gen}}$. We repeat all the experiments at least 10 times and average the results over the runs. Since our main objective is to showcase how the model-based continuation ripples out to the behavior of the unfolded network, we only employ DECONET as a baseline and leave other DUNs as future work, especially since [deconet] has demonstrated that DECONET outperforms two other state-of-the-art DUNs. For the loss landscapes, we follow the standard framework of [lossvis] and perturb the learned $W$ in two random directions, with scalars $Alpha$ and $Beta$ representing the extent of the perturbation in each direction.
@@
Results & discussion: We examine the reconstruction and generalization performance of 10- and 15-layer C-DEC, with $W\in\mathbb{R}^{10n\times n}$ and varying number of continuation steps, on both datasets, and illustrate the findings in Fig. 1. From a model-based viewpoint, the decays in test errors seem reasonable: as the number of continuation steps (and layers) increases, the solver’s reconstruction ability is improved [tfocs]. This behavior conforms with our intuition presented in Sec. II. Interestingly, we find that the number of continuation steps is mildly reversely proportional to the generalization error. We surmise that this is due to the nested architecture of C-DEC compared to DECONET, although further mathematical exploration is needed. Additionally, we compare C-DEC and DECONET, both with $W\in\mathbb{R}^{50n\times n}$, for 10 and 50 layers, on all datasets, and present the findings in Table I. Both the test and generalization errors are always lower for our proposed DUN, consistently for both datasets. Based on the results, we obtain a two-fold insight: a) we confirm our model-based intuition that continuation boosts the solver’s performance, even in the unfolding regime b) we justify our choice of the log-cosh loss as an adequate metric for training and evaluating the proposed DUN, compared to MSE, which is deployed for DECONET. Overall, results highlight that continuation is a favorable component of C-DEC’s architecture, with the log-cosh loss being a better choice over standard metrics, e.g., the MSE, for measuring DUNs’ performance.
Finally, we examine how continuation ripples out to C-DEC’s loss landscapes. We compare a 5-layer C-DEC without continuation – which is architecturally equivalent to DECONET – and a 5-layer C-DEC with 5 steps of continuation. We examine both DUNs with $W\in\mathbb{R}^{10n\times n}$ and $W\in\mathbb{R}^{50n\times n}$, and report the results in Fig. 2. The illustrations demonstrate a very intriguing phenomenon. Firstly, we observe that for both datasets, DECONET (equivalent to C-DEC with 1-step continuation) suffers from a highly nonsmooth loss landscape, with many peaks and valleys. On the other hand, C-DEC with 5 steps of continuation enjoys a much smoother loss landscape, with the effect being highly noticeable for the CIFAR10 images, as depicted in Fig. 2b. Then, for the MNIST dataset illustrated in Fig. 2a, we notice that overparameterization also smooths out the landscape, as $N$ increases from $10n$ to $50n$. Although this effect is negligible for DECONET, it is more noticeable for C-DEC, showcasing the potential of coupling continuation with overparameterization in the context of DUNs. All in all, results indicate that continuation (and overparameterization) has a strong positive effect on the optimization process of DUNs, and could spark an interesting mathematical theory regarding DUNs’ generalization.